OpenStudy (anonymous):

The width of a rectangle is (3x − 3) cm. The length of the rectangle is twice the width. Find the area of the rectangle in terms of the variable x.

OpenStudy (mathteacher1729):

Area = Length times Width. Width = (3x-3) factor out a 3 Width = 3(x-1) Length = 2*3(x-1) Multiply 2*3 = 6 Length = 6(x-1) Length * Width = 3(x-1)*6(x-1) multiply the 3*6 Length * Width = 18(x-1)(x-1) Can you expand (x-1)(x-1) an then distribute the 18?